OpenStudy (migitmack):

Julio joined a golf club two years ago. He pays an annual membership fee of $1,150 and a greens fee of $25 each time he plays a game of golf. The function below can be used to calculate his total yearly golfing fee, f (g), in dollars. f (g) = 1150 + 25g, where g represents the number of times he played golf during the year. Last year he paid $2,325 as a total golfing fee. For how many games did he pay a greens fee?

OpenStudy (ganpat):

2 yr , he is playing.. and let x = be no of games played.. (1150) + 25 (x) = 2325.. 25x = 2325 - 1150 25x = 1175 x = 47... number of gamess..